Calculus II, Spring 2016 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 14

1. Section 15.8: Problems 5, 7, 9, 15, 23.

2. Chap 15, additional and advanced exercises: Problems 11, 12.

3. Section 16.1: Problems 15, 23, 25, 29.

4. Section 16.2: Problems 19, 23, 25, 27, 29, 35, 47.

See equation (5), (6) for definition of flow, circulation and flux.